# Preprint submitted to Mathematical Geosciences

### v1

@@ 3.1 Diffusion model (DM)
As noted in Di Federico and Durlofsky (2025), LDMs require a compromise between dimension reduction ratio $f$ and generation quality, and in this study, we use $f=8$ (overall $f^{3}=512$). The same dataset split, optimizer, learning rate, and batch size applied for the training of the VAE is used here. For the U-net, training requires approximately 8 h (300 epochs) on the same NVIDIA V100 GPU. We adopt a DDIM sampling and a linear scheduler for $T=20$ steps, with $\alpha_{1}=0.9985$ and $\alpha_{T}=0.9805$.
The implementation developed for this study is based on the libraries diffusers and monai (see the “Code availability” section for additional details). The detailed architectures used in this work are provided in Table 3 for the VAE and Table 4 for the U-net (in Appendix 7).
After both the VAE and the U-net are trained, a new realization ${\mathbf{m}}_{0}^{\text{LDM}}$ can be generated by iteratively denoising a random sample $\bm{\xi}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n_{c}})$. A visual representation of this process is given in Figure 3a. The generic latent variable $\bm{\xi}_{t}$ is gradually denoised using the predicted noise from the U-net, $\bm{\epsilon}_{\theta}(\bm{\xi}_{t},t)$, together with the step $t$ and parameter $\bar{\alpha}_{t}$ from the scheduler, through Eq. 8. The denoised latent variable $\bm{\xi}_{t-1}$ is obtained at each step. When the last step is reached, i.e., $t=1$, the latent variable is decoded into the generated geomodel through application of $\mathbf{m}_{0}^{\text{LDM}}=\mathcal{D}(\bm{\xi}_{0})$. Figure 3b shows an example of geomodels generated by decoding the fully denoised latent variable $\bm{\xi}_{0}$, for different numbers of total DDIM steps ($T$), starting from the same input $\bm{\xi}_{T}$. It can be observed that, for $T>10$, the overall quality remains stable, and extending the denoising process only provides a cleaner, more noise-free geomodel. The images here represent the generated geomodel for different $T$ values. This is for illustration purposes only because, in the actual implementation, the latent variable is decoded only at the final step.
@@ 4 Geomodel generation and flow simulations with 3D-LDM
We now employ the 3D-LDM to generate realizations of the three-facies channelized system. To evaluate the similarity between LDM generated models and reference models, we utilize both qualitative visual assessments and quantitative metrics. The latter include spatial statistics and flow simulation results. Though the main purpose of our LDM procedure is to provide a parameterized representation that can be used for history matching, the approach also leads to significant computational savings in the model generation step itself. Specifically, using the same NVIDIA V100 GPU, only 2 minutes are required to generate 1000, $\times128$$\times$32 realizations, compared to several hours using the Petrel OBM tool on an Intel Core i7-7700 CPU.
@@ 4.1 Qualitative assessments
The scenario parameters for Petrel-generated samples are prescribed as OBM inputs, so they are immediately available. This is not the case for 3D-LDM realizations, however, since the only input is a randomly generated $\bm{\xi}_{T}$ field. Thus, if scenario parameters are of interest, these must be determined from the LDM-generated geomodels. Mud fraction can be directly computed as the fraction of 0-facies cells. To efficiently determine the channel orientation and channel width for LDM-generated realizations, we train a separate convolutional neural network (CNN) to provide these quantities given the input geomodel. This CNN is trained on the same dataset as was used for the 3D-LDM training. To simplify this procedure, only layers $z=1$ and $z=16$ are used, as we have found these to be representative of the overall geomodel scenario parameters. The average testing RMSE with this approach is about 2.3 and 0.15 grid blocks for channel orientation and width. These errors are comparable to the variability introduced by the OBM drift discussed in Section 2, so we believe this treatment is sufficient for current (reporting) purposes.
The scenario parameters computed in this way for the LDM-generated geomodels are provided in Figure 4. It is evident that the geomodels correspond to a variety of scenario parameters – high and low mud fractions, different orientations, wide and thin channels – consistent with the training set. This is an important observation as it suggests that the full range of models comprising the training set can indeed be generated. Another model characteristic worth assessing is the level of accuracy in hard data conditioning. We observe that hard data at conditioning locations are honored with an accuracy of 98% over the full set of LDM generated models. This could be improved by increasing $\lambda_{h}$ in Eq. 4, though that could negatively impact other geomodel features.
(a) Mud fraction: 0.74
@@
Figure 4: Randomly selected 3D-LDM generated realizations that span different geological scenarios. Scenario parameters are reported for each geomodel.
@@
Figure 6 highlights the impact of the perceptual loss term, $\lambda_{\text{perc}}L_{\text{perc}}$ in Eq. 4, in the VAE training. This loss term significantly reduces (essentially eliminates) the occurrence of unrealistic features in the decoded geomodels such as broken channel bodies, incorrect facies ordering, or isolated levee cells. Such unwanted features are visible in the magnified (circular) views in the areal and vertical cross section planes in Figure 6a, where the perceptual loss is not included. There we see some “speckling” where levee facies appear within channel facies. This issue is not observed in Figure 6b, where the perceptual loss is included. Although a one-to-one comparison is not possible as the samples are generated by LDMs with different loss functions and hence different parameters, the model quality shown in Figure 6b is consistently observed. This demonstrates that the use of perceptual loss is indeed beneficial for our multiscenario modeling.
@@ 4.2 Quantitative assessments
Variograms are commonly used in geostatistics to quantify two-point spatial correlations. These fully characterize multi-Gaussian fields, and provide useful (though incomplete) metrics for more general multipoint geostatistical models, such as those considered here. For a 2D slice $i$ along axis $d$ of a geomodel, $\mathbf{m}_{0,i}^{(d)}$, the variogram (or “semi-variance”) $\gamma(h_{\theta})$ is a function of the lag distance $h_{\theta}$ in the direction aligned with angle $\theta$ (Pyrcz and Deutsch, 2014), i.e.,
@@
We now evaluate the performance of our 3D-LDM parameterization in a flow simulation setting. Flow simulations are performed for a water-flooding problem involving a two-phase immiscible oil-water system. The setup entails three water injection wells and six production wells arranged in a line-drive pattern, as shown in Figure 1a. Grid blocks are of physical dimensions 65 ft $\times$ 65 ft $\times$ 6.5 ft (in the $x$, $y$ and $z$ directions, respectively). The reservoir is initially at a pressure of 4500 psi (at a depth of 4160 ft). Initial oil and water saturations are 0.85 and 0.15, and their viscosities are 5 cP and 1 cP, respectively, at reference conditions. The relative permeability curves are the same as those used in Di Federico and Durlofsky (2025). Porosity and permeability are constant for each facies. The specific values are 0.2 and 1000 mD for channels, 0.15 and 200 mD for levee, and 0.05 and 25 mD for mud. Wells are all bottom-hole pressure (BHP) controlled, with injectors operating at 4700 psi and producers at 4300 psi. The wells are perforated only at grid blocks corresponding to channel facies (see Figure 1a). The simulation time frame is 1500 days. All runs are performed using the tNavigator simulator (Rock Flow Dynamics, 2024). Results for flow statistics are shown in Figure 9. For both the Petrel and 3D-LDM ensembles, we present the 10th, 50th and 90th percentile flow response (P10, P50 and P90) over the full set of runs at each time step. These results are at the field level, i.e., summed over all three injectors (Figure 9a) or all six producers (Figure 9b,c). These flow statistics display very close agreement – the only visible discrepancies are slight deviations in the P90 water injection (Figure 9a) and water production (Figure 9c) curves. We note that consistency is also observed at the well level (these results are not shown for conciseness). The agreement highlighted in Figure 9 is important because it suggests that the range of spatial features that impact flow are indeed captured in the LDM generated models.
@@
The results in this section demonstrate that our 3D-LDM procedure is accurate for the range of scenario parameters considered (see Table 1).
For somewhat larger parameter ranges (e.g., mud fraction ranging from 0.6 to 0.9, average channel orientation from 0 to 60, average channel width from 3 to 7 grid blocks), the 3D-LDM did not perform as well. For such cases, the generated models still appeared geologically realistic (visually), though they did not span the full range of the training models. Specifically, we observed incomplete coverage of the ranges in the generated geomodels for ensemble metrics such as those shown in Figures 7 and 8.
@@ 5 Latent space history matching with uncertain geological scenario
In this section, we apply the 3D-LDM within the context of history matching. The flow problem and simulation setup are as described in Section 4.2. For discrete facies geomodels such as those considered in this study, the goal of the data assimilation procedure is to assign a facies type to each grid block.
In the multiscenario case, this process is more challenging because the latent variables must characterize (indirectly) the specific scenario parameters, which in this case correspond to mud fraction, channel orientation, and channel width. For the purposes of history matching, the training dataset (or a statistically representative subset) defines the prior.
@@ 5.1 History matching cases and procedure
Mud fraction 0.75 0.85 0.81
Channel orientation (avg. angle) 50 59 32
Channel width (avg. # of blocks) 5.9 4.3 4.6
History matching updates are performed in the LDM latent space, such that the variables to calibrate correspond to $\bm{\xi}_{T}\in\mathbb{R}^{n_{c}}$. This reduces the number of history matching variables from $N_{c}=528,288$ (i.e., with no parameterization) to $n_{c}=N_{c}/f^{3}=1024$. At every iteration, the 3D-LDM parameterization generates the facies model corresponding to the updated $\bm{\xi}_{T}$, $\mathbf{m}_{0}^{\text{LDM}}(\bm{\xi}_{T})$, through the procedure shown in Figure 3(a). Grid blocks are then assigned the appropriate (facies-based) porosity and permeability values used in Section 4.2. Flow simulations are performed in the geomodel space. Note that, as demonstrated for a 2D case in Di Federico and Durlofsky (2025), it is straightforward to include the facies properties (porosity and permeability) in the set of unknowns to be determined during history matching, though this is not considered here.
@@
The widely used ensemble smoother with multiple data assimilation (ESMDA), developed by Emerick and Reynolds (2013), is applied in this work. We apply ESMDA to the latent variables, and do not include any specialized treatments such as localization. In ESMDA, an ensemble of $N_{e}$ $\bm{\xi}_{T}$ fields is updated at every iteration $u$ based on the mismatch between observed data $\mathbf{d}_{\text{obs}}^{*}$ and simulated data $\mathbf{d}$, obtained by performing flow simulation for each geomodel in the ensemble. The update equation is
@@
where $C_{\xi d}\in\mathbb{R}^{n_{c}\times N_{d}}$ is the cross-covariance matrix between $\bm{\xi}_{T}$ and $\mathbf{d}$, and $C_{dd}\in\mathbb{R}^{N_{d}\times N_{d}}$ is the autocovariance of $\mathbf{d}$. The value $\alpha$ is an inflation coefficient which must satisfy $\sum_{i=1}^{N_{a}}{\alpha_{i}}^{-1}=1$. Here we set $N_{e}=200$, $N_{a}=10$, and $\alpha_{i}$ values to [57.017, 35.0, 25.0, 20.0, 18.0, 15.0, 12.0, 8.0, 5.0, 3.0], as suggested by Emerick and Reynolds (2013).
The discrete, non-Gaussian nature of the facies models considered here would present challenges for ESMDA applied directly in the geomodel space (i.e., with no parameterization). For example, directly updating individual grid blocks over data assimilation steps may not preserve the continuity or geological character of the geological structures (though treatments such as localization can be used to mitigate this effect). The use of the 3D-LDM parameterization addresses this concern since the latent variable $\bm{\xi}_{T}$ is normally distributed, consistent with ESMDA assumptions, and the 3D-LDM mapping maintains geological realism after latent variable updates.

### v2

@@ 3.1 Diffusion model (DM)
As noted in DIFEDERICO2025105755, LDMs require a compromise between dimension reduction ratio $f$ and generation quality. In this study, we performed numerical experimentation with $f=4$, 8 and 16. We found the use of $f=8$ (overall $f^{3}=512$) to provide the best balance, i.e., a sufficiently low-dimension latent space and fast geomodel generation, while also achieving minimal information loss from encoding. Thus $f=8$ is used in all results in this paper. The same dataset split, optimizer, learning rate, and batch size applied for the training of the VAE are used here. For the U-net, training requires approximately 8 h (300 epochs) on the same NVIDIA A100 GPU. We adopt a DDIM sampling and a linear scheduler for $T=20$ steps, with $\alpha_{1}=0.9985$ and $\alpha_{T}=0.9805$.
The implementation developed for this study is based on the libraries diffusers and monai (see the “Code availability” section for additional details). The detailed architectures used in this work are provided in Table 3 for the VAE and Table 4 for the U-net (in Appendix 3D-LDM model architecture).
After both the VAE and the U-net are trained, a new realization ${\mathbf{m}}_{0}^{\text{LDM}}$ can be generated by iteratively denoising a random sample $\bm{\xi}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n_{c}})$. A visual representation of this process is given in Figure 3a. The generic latent variable $\bm{\xi}_{t}$ is gradually denoised using the predicted noise from the U-net, $\bm{\epsilon}_{\theta}(\bm{\xi}_{t},t)$, together with the step $t$ and parameter $\bar{\alpha}_{t}$ from the scheduler, through Eq. 8. The denoised latent variable $\bm{\xi}_{t-1}$ is obtained at each step. When the last step is reached, i.e., $t=1$, the latent variable is decoded into the generated geomodel through application of $\mathbf{m}_{0}^{\text{LDM}}=\mathcal{D}(\bm{\xi}_{0})$. Pseudocode for the generation (inference) procedure is provided in Algorithm 3 in Appendix 3D-LDM algorithms.
Figure 3b shows an example of geomodels generated by decoding the fully denoised latent variable $\bm{\xi}_{0}$, for different numbers of total DDIM steps ($T$), starting from the same input $\bm{\xi}_{T}$. It can be observed that, for $T>10$, the overall quality remains stable, and extending the denoising process only provides a cleaner, more noise-free geomodel. The images here represent the generated geomodel for different $T$ values. This is for illustration purposes only because, in the actual implementation, the latent variable is decoded only at the final step.
@@ 4 Geomodel generation and flow simulations with 3D-LDM
We now employ the 3D-LDM to generate realizations of the three-facies channelized system. To evaluate the similarity between LDM generated models and reference models, we utilize both qualitative visual assessments and quantitative metrics. The latter include spatial statistics and flow simulation results. Though the main purpose of our LDM procedure is to provide a parameterized representation that can be used for history matching, the approach also leads to significant computational savings in the model generation step itself. Specifically, using the same NVIDIA A100 GPU, only 2 minutes are required to generate 1000, 128 $\times$ 128 $\times$ 32 realizations, compared to several hours using the Petrel OBM tool on an Intel Core i7-7700 CPU.
@@ 4.1 Qualitative assessments
The scenario parameters $\mathbf{s}=[f_{m},\theta_{\rm ch},w_{\rm ch}]$ for Petrel-generated samples are prescribed as OBM inputs, so they are immediately available. This is not the case for 3D-LDM realizations, however, since the only input is a randomly generated $\bm{\xi}_{T}$ field. Thus, if scenario parameters are of interest, these must be determined from the LDM-generated geomodels. Mud fraction can be directly computed as the fraction of 0-facies cells. To efficiently determine the channel orientation and channel width for LDM-generated realizations, we train a separate convolutional neural network (CNN) to provide these quantities given the input facies model $\mathbf{r}$. We denote this estimation of $\theta_{\rm ch}$ and $w_{\rm ch}$ as $\tilde{\mathbf{s}}=\text{CNN}(\mathbf{r})$. This CNN is trained on the same dataset as was used for the 3D-LDM training. To simplify this procedure, only layers $z=1$ and $z=16$ are used, as we have found these to be representative of the overall geomodel scenario parameters. The average testing RMSE with this approach is about 2.3 and 0.15 grid blocks for channel orientation and width. These errors are comparable to the variability introduced by the OBM drift discussed in Section 2, so we believe this treatment is sufficient for current (reporting) purposes.
The scenario parameters computed in this way for the LDM-generated geomodels are provided in Figure 4. It is evident that the geomodels correspond to a variety of scenario parameters – high and low mud fractions, different orientations, wide and thin channels – consistent with the training set. This is an important observation as it suggests that the full range of models comprising the training set can indeed be generated. Another model characteristic worth assessing is the level of accuracy in hard data conditioning. We observe that hard data at conditioning locations are honored with an accuracy of 98% over the full set of LDM generated models. This could be improved by increasing $\lambda_{\text{h}}$ in Eq. 4, though that could negatively impact other geomodel features.
Mud fraction: 0.74
@@
Figure 4: Randomly selected 3D-LDM generated realizations corresponding to different scenario parameters. Scenario parameters are reported for each geomodel.
@@
Figure 6 highlights the impact of the perceptual loss term, $\lambda_{\text{perc}}L_{\text{perc}}$ in Eq. 4, in the VAE training. This loss term significantly reduces (essentially eliminates) the occurrence of unrealistic features in the decoded geomodels such as broken channel bodies, incorrect facies ordering, or isolated levee cells. Such unwanted features are visible in the magnified (circular) views in the areal and vertical cross section planes in Figure 6a, where the perceptual loss is not included. There we see some “speckling” where levee facies appear within channel facies. This issue is not observed in Figure 6b, where the perceptual loss is included. Although a one-to-one comparison is not possible as the samples are generated by LDMs with different loss functions and hence different parameters, the model quality shown in Figure 6b is consistently observed. This demonstrates that the use of perceptual loss is indeed beneficial for our variable scenario modeling.
@@ 4.2 Quantitative assessments
Variograms are commonly used in geostatistics to quantify two-point spatial correlations. These fully characterize multi-Gaussian fields, and provide useful (though incomplete) metrics for more general multipoint geostatistical models, such as those considered here. For a 2D slice $i$ along axis $d$ of a geomodel, $\mathbf{m}_{0,i}^{(d)}$, the variogram (or “semi-variance”) $\gamma(h_{\theta})$ is a function of the lag distance $h_{\theta}$ in the direction aligned with angle $\theta$ (pyrcz2014geostatistical), i.e.,
@@
We now evaluate the performance of our 3D-LDM parameterization in a flow simulation setting. Flow simulations are performed for a water-flooding problem involving a two-phase immiscible oil-water system. The setup entails three water injection wells and six production wells arranged in a line-drive pattern, as shown in Figure 1a. Grid blocks are of physical dimensions 65 ft $\times$ 65 ft $\times$ 6.5 ft (in the $x$, $y$ and $z$ directions, respectively). The reservoir is initially at a pressure of 4500 psi (at a depth of 4160 ft). Initial oil and water saturations are 0.85 and 0.15, and their viscosities are 5 cP and 1 cP, respectively, at reference conditions. The relative permeability curves are the same as those used in DIFEDERICO2025105755. Porosity and permeability are constant for each facies. The specific values are 0.2 and 1000 mD for channels, 0.15 and 200 mD for levee, and 0.05 and 25 mD for mud. Wells are all bottom-hole pressure (BHP) controlled, with injectors operating at 4700 psi and producers at 4300 psi. The wells are perforated only at grid blocks corresponding to channel facies (see Figure 1a). The simulation time frame is 1500 days. All runs are performed using the tNavigator simulator (rock_flow_dynamics). Using 16 CPU cores, a single simulation run requires about 10 minutes.
Results for flow statistics are shown in Figure 9. For both the Petrel and 3D-LDM ensembles, we present the 10th, 50th and 90th percentile flow response (P10, P50 and P90) over the full set of runs at each time step. These results are at the field level, i.e., summed over all three injectors (Figure 9a) or all six producers (Figure 9b,c). These flow statistics display very close agreement – the only visible discrepancies are slight deviations in the P90 water injection (Figure 9a) and water production (Figure 9c) curves. We note that consistency is also observed at the well level (these results are not shown for conciseness). The agreement highlighted in Figure 9 is important because it suggests that the range of spatial features that impact flow are indeed captured in the LDM generated models.
@@
The results in this section demonstrate that our 3D-LDM procedure is accurate for the range of scenario parameters considered (see Table 1). For somewhat larger parameter ranges (e.g., $f_{m}$ ranging from 0.6 to 0.9, $\theta_{\rm ch}$ from 0 to 60, $w_{\rm ch}$ from 3 to 7 grid blocks), the 3D-LDM did not perform as well. For such cases, the generated models still appeared geologically realistic (visually), though they did not span the full range of the training models. Specifically, we observed incomplete coverage of the ranges in the generated geomodels for ensemble metrics such as those shown in Figures 7 and 8.
@@ 5 Latent space history matching with uncertain geological scenario
In this section, we apply the 3D-LDM within the context of history matching. The flow problem and simulation setup are as described in Section 4.2. For discrete facies geomodels such as those considered in this study, the goal of the data assimilation procedure is to assign a facies type to each grid block. In the hierarchical uncertainty case, this process is more challenging because the latent variables must characterize (indirectly) the specific scenario parameters, which in this case correspond to mud fraction, channel orientation, and channel width.
For the purposes of history matching, the training dataset (or a statistically representative subset, as assumed in Section 4.2) defines the prior. In other words, the history matching method applied here assumes the priors have been “validated” and that the true model and data lie within the spaces defined by the training dataset. If this is not the case, the priors should be modified accordingly, or a treatment for model error that accounts for this discrepancy should be incorporated. Either or both of these steps must be performed before the history matching procedure described in Section 5.1 is applied.
@@ 5.1 History matching cases and procedure
Mud fraction, $f_{m}$ 0.75 0.85 0.81
Channel orientation (avg. angle), $\theta_{\rm ch}$ 50 59 32
Channel width (avg. # of blocks), $w_{\rm ch}$ 5.9 4.3 4.6
History matching updates are performed in the LDM latent space, such that the variables to calibrate correspond to $\bm{\xi}_{T}\in\mathbb{R}^{n_{c}}$. This reduces the number of history matching variables from $N_{c}=524,288$ (i.e., with no parameterization) to $n_{c}=N_{c}/f^{3}=1024$. At every iteration, the 3D-LDM parameterization generates the facies model corresponding to the updated $\bm{\xi}_{T}$, $\mathbf{m}_{0}^{\text{LDM}}(\bm{\xi}_{T})$, through the procedure shown in Figure 3(a). Grid blocks are then assigned the appropriate (facies-based) porosity and permeability values used in Section 4.2. Flow simulations are performed in the geomodel space. Note that, as demonstrated for a 2D case in DIFEDERICO2025105755, it is straightforward to include the facies properties (porosity and permeability) in the set of unknowns to be determined during history matching, though this is not considered here.
@@
Before describing the specific history matching procedure applied in this study, we discuss the hierarchical data assimilation problem that is being treated. Recall that we can express the geomodel as $\mathbf{m}=\mathbf{m}(\mathbf{s},\mathbf{r})$, where $\mathbf{s}$ and $\mathbf{r}$ refer to scenario parameters and cell-by-cell facies values, respectively. With this representation, and consistent with, e.g., roininen2016hyperpriorsmaternfieldsapplications and TENG2025104961, the joint posterior density in a Bayesian framework, denoted $p(\mathbf{s},\mathbf{r}|\mathbf{d}^{*}_{\text{obs}})$, is given by
$p(\mathbf{s},\mathbf{r}|\mathbf{d}^{*}_{\text{obs}})=\frac{p(\mathbf{d}^{*}_{\text{obs}}|\mathbf{s},\mathbf{r})\,p(\mathbf{s},\mathbf{r})}{p(\mathbf{d}^{*}_{\text{obs}})}.$ (11)
Here $p(\mathbf{d}^{*}_{\text{obs}}|\mathbf{s},\mathbf{r})$ is the likelihood function, $p(\mathbf{s},\mathbf{r})$ is the prior density, and $p(\mathbf{d}^{*}_{\text{obs}})$ is the evidence, which acts as a normalization constant. Following (TENG2025104961), the joint posterior can be decomposed into the marginal posterior of $\mathbf{s}$ and the conditional posterior of $\mathbf{r}$ given $\mathbf{s}$
$p(\mathbf{s},\mathbf{r}|\mathbf{d}^{*}_{\text{obs}})=p(\mathbf{s}|\mathbf{d}^{*}_{\text{obs}})\,p(\mathbf{r}|\mathbf{s},\mathbf{d}^{*}_{\text{obs}}).$ (12)
In turn, $p(\mathbf{s}|\mathbf{d}^{*}_{\text{obs}})$ is given by Bayes’ rule as
$p(\mathbf{s}|\mathbf{d}^{*}_{\text{obs}})=\frac{p(\mathbf{d}^{*}_{\text{obs}}|\mathbf{s})\,p(\mathbf{s})}{p(\mathbf{d}^{*}_{\text{obs}})},$ (13)
where $p(\mathbf{s})$ is the prior of the scenario parameters. The likelihood in Eq. 13 is difficult to compute because it entails marginalization over $\mathbf{r}$, i.e.,
$p(\mathbf{d}^{*}_{\text{obs}}|\mathbf{s})=\int p(\mathbf{d}^{*}_{\text{obs}}|\mathbf{r},\mathbf{s})\,p(\mathbf{r}|\mathbf{s})\,d\mathbf{r}.$ (14)
In our parameterization, the variables $(\mathbf{s},\mathbf{r})$ are replaced with the latent variable $\bm{\xi}_{T}$. The data assimilation problem is thus simplified considerably, as we now only need to estimate the posterior distribution $p(\bm{\xi}_{T}|\mathbf{d}^{*}_{\text{obs}})$ instead of $p(\mathbf{s},\mathbf{r}|\mathbf{d}^{*}_{\text{obs}})$. In addition, in contrast to s and r, $\bm{\xi}_{T}$ is normally distributed.
We apply the widely used ensemble smoother with multiple data assimilation (ESMDA) procedure, developed by Emerick and Reynolds EMERICK20133, for history matching, i.e., to sample the posterior $p(\bm{\xi}_{T}|\mathbf{d}^{*}_{\text{obs}})$. ESMDA operates directly on the latent variables. No specialized treatments such as localization are used. Because the relationship between $\bm{\xi}_{T}$ and spatial locations is indirect, distance-based localization could be challenging to apply. Correlation-based localization LACERDA2019690, which leverages the sample cross-correlation between latent variables and observed data, could be considered. This did not appear to be necessary for the examples in this paper.
In ESMDA, an ensemble of $\bm{\xi}_{T}$ fields is updated (superscript $u$) at every iteration based on the mismatch between observed data $\mathbf{d}_{\text{obs}}^{*}$ and simulated data $\mathbf{d}$, obtained by performing flow simulation for each geomodel in the ensemble. The update equation is
@@
where $N_{e}$ is the number of ensemble members, $C_{\xi d}\in\mathbb{R}^{n_{c}\times N_{d}}$ is the cross-covariance matrix between $\bm{\xi}_{T}$ and $\mathbf{d}$, and $C_{dd}\in\mathbb{R}^{N_{d}\times N_{d}}$ is the autocovariance of $\mathbf{d}$. The value $\alpha$ is an inflation coefficient, which must satisfy $\sum_{i=1}^{N_{a}}{\alpha_{i}}^{-1}=1$. Here we set $N_{e}=200$, $N_{a}=10$, and $\alpha_{i}$ values to [57.017, 35.0, 25.0, 20.0, 18.0, 15.0, 12.0, 8.0, 5.0, 3.0], as suggested by EMERICK20133. Since the ensemble members are independent, simulations can be performed in parallel. We simulate these models in batches of 20, which results in an elapsed time of approximately 1.5 hours per ESMDA iteration, or about 15 hours total for a full history matching run.
The discrete, non-Gaussian nature of the facies models considered here would present challenges for ESMDA applied directly in the geomodel space (i.e., with no parameterization). For example, directly updating individual grid blocks over data assimilation steps may not preserve the continuity or geological character of the geological structures (though localization can be used to mitigate this effect). The use of the 3D-LDM parameterization addresses this concern since the latent variable $\bm{\xi}_{T}$ is normally distributed, consistent with ESMDA assumptions, and the 3D-LDM mapping maintains geological realism after latent variable updates.
On the other hand, it should be noted that the parameterization introduces an additional nonlinear mapping step between the geological model and simulated data. For discrete-facies models, this can result in a discontinuity between scenario parameters and grid-block facies values. As a consequence, $C_{\xi d}$ may only approximate the relationship between $\bm{\xi}_{T}$ and $\mathbf{d}$. Nevertheless, the results presented below confirm the effectiveness of the 3D-LDM-based ESMDA procedure for our cases. This effective performance may be related to the fact that the data used here are well rates, which involve summation over multiple geological layers. Thus, local changes in facies type may have a relatively small impact on d (recall also that we condition to hard data at well locations). In future work, we plan to evaluate the performance of 3D-LDM with other history matching methods, such as Markov chain Monte Carlo.